# Sim2Real Transfer for Audio-Visual Navigation with Frequency-Adaptive Acoustic Field Prediction

### v1

@@ Abstract
Sim2real transfer has received increasing attention lately due to the success of learning robotic tasks in simulation end-to-end. While there has been a lot of progress in transferring vision-based navigation policies, the existing sim2real strategy for audio-visual navigation performs data augmentation empirically without measuring the acoustic gap. The sound differs from light in that it spans across much wider frequencies and thus requires a different solution for sim2real. We propose the first treatment of sim2real for audio-visual navigation by disentangling it into acoustic field prediction (AFP) and waypoint navigation. We first validate our design choice in the SoundSpaces simulator and show improvement on the Continuous AudioGoal navigation benchmark. We then collect real-world data to measure the spectral difference between the simulation and the real world by training AFP models that only take a specific frequency subband as input. We further propose a frequency-adaptive strategy that intelligently selects the best frequency band for prediction based on both the measured spectral difference and the energy distribution of the received audio, which improves the performance on the real data. Lastly, we build a real robot platform and show that the transferred policy can successfully navigate to sounding objects. This work demonstrates the potential of building intelligent agents that can see, hear, and act entirely from simulation, and transferring them to the real world.
@@ I Introduction
Navigation has been extensively studied in the robotics community and has been traditionally approached with Simultaneous Localization and Mapping (SLAM) [1] with Lidar sensors. This approach however is limited to geometry planning and does not reason well about the semantics of the scene. In addition, building robots and reproducing experiments are expensive for real robots.
To address these issues, recent years have witnessed research approaching the navigation problem from a vision-centric perspective, i.e., the robot mainly uses visual sensors to perceive the scene [2]. This approach has demonstrated a lot of success in photorealistic real-scanned environments, which allow fast experimentation and replicable research. Various forms of tasks have been proposed in this domain including using egocentric vision to travel to a designated point in an unfamiliar environment [3, 4, 5], search for a specified object [6, 7], or explore and map a new space [8, 9, 10]. Some other works further explore expanding the sensory suite of the navigating agent to include hearing as well. In particular, the AudioGoal task [11, 12] requires an agent to navigate to a sounding target (e.g., a ringing phone) using audio for directional and distance cues while using vision to avoid obstacles in the unmapped environment.
With the success of these learning-based navigation systems in photorealistic simulation environments, some work explores transferring the learned policy to the real world by bridging the gap between the simulation and the real world [13, 14, 15, 16]. Recent work [17] does sim2real transfer for audio-visual navigation with data augmentation however without further investigating the acoustic gap. The sound differs from light in that it spans a wide range of frequencies, which is one of the main barriers to sim2real transfer. In this work, we perform a systematic evaluation of the acoustic gap and propose a solution to bridge that gap.
@@
State-of-the-art approaches in audio-visual navigation rely on reinforcement learning to train the navigation policy end-to-end [11, 18], which is not only hard to interpret but also impractical to generalize to the real world directly due to various sim2real gaps. Recent visual navigation work has shown success in sim2real transfer with hierarchical models [14, 8], which typically consist of a high-level path planner and low-level motion planner. This design helps abstract away some of the low-level physical discrepancies.
@@
Inspired by such methods, we design a modular approach to ease the transfer from the simulation to the real world. To achieve this, we confront a key question: what is the proper high-level planning task that can survive sim2real transfer for audio-visual navigation? To this end, we propose a novel prediction task: acoustic field prediction—predicting the local sound pressure field around the agent. The gradient of this field reflects the direction of the sound). Measuring acoustic fields is expensive in the real world since it requires simultaneously capturing the sound pressure of all points in the field due to the dynamic nature of sound. However, they are free to compute in simulation. We first build an audio-visual model as the acoustic field predictor (AFP) and curate a large-scale acoustic field dataset on SoundSpaces 2.0 [19], the state-of-the-art audio-visual simulation platform. We show that this approach outperforms existing methods on the Continuous AudioGoal navigation benchmark.
@@
After validating the proposed approach in simulation, we then investigate where acoustic discrepancy arises. It is known that ray-tracing-based acoustic simulation algorithms introduce more errors with lower frequencies due to wave effects [20]. Given this observation, we focus on evaluating how the sim2real error changes as a function of frequencies. We first collect real acoustic field data with the source sound being white noise, whose audio energy uniformly spans across all frequencies. We then train acoustic field prediction models that only take the sub-frequency band of the input audio and test it on the real white noise data. By computing the errors across multiple samples, we show that the errors do not strictly go down as the frequency goes higher, and using the best frequency band yields errors smaller than using all frequencies for the white noise sound.
@@
However, simply taking the best frequency band doesn’t work for all sounds since different sounds have different spectral distributions. To address this issue and make the model aware of the spectral difference, we propose a novel frequency-adaptive prediction strategy, that intelligently selects the best frequency sub-band based on measured errors as well as the received spectral distribution to predict the acoustic field. To validate this approach, we collect more acoustic field data with various sounds and show that the frequency-adaptive model leads to the lowest error on the real data compared to other strategies.
@@
In summary, we propose a novel acoustic field prediction approach that learns to navigate without interaction with the environment. This approach improves the SOTA methods on the challenging Continuous AudioGoal navigation benchmark. We perform a systematic evaluation of the sim2real and propose a frequency-adaptive strategy as the treatment for sim2real. We show this strategy works on both collected real data as well as on our robot platform. To the best of our knowledge, this is the first work to investigate and propose a principal solution to the sim2real transfer problem for audio-visual navigation.
@@ II-A Embodied Navigation
To navigate autonomously, traditionally a robot builds a map via 3D reconstruction (i.e., SLAM) and then plans a path using the map [21]. Recent works have developed navigation policies that make navigation decisions in a previously unmapped environment from egocentric observations directly without relying on mapping [4, 7, 22]. Some recent efforts have developed audio-visual simulation platforms [11, 19, 23] that enable embodied agents to both see and hear. AudioGoal navigation [11, 18] is also introduced where the agent is tasked to navigate to the source of sound in an unknown environment. The state-of-the-art audio-visual navigation models train policies with reinforcement learning (RL) that requires millions of samples. Inspired by recent work in learning potential functions for interaction-free navigation [8], we propose to predict acoustic fields for interaction-free audio-visual navigation.
@@ II-C Sim2real Transfer
Benefiting from recent large-scale datasets of real-world 3D scans [36, 37] and supporting simulators  [5, 38], recent works have shown success in enabling embodied agents in the simulation. Transferring the model trained in simulation to the real world is thus of great interest. The mostly widely used approaches are domain randomization [39, 40], system identification [16, 41], and transfer learning and domain adaptation [42]. Most sim2real research studies transferring a policy from simulation to the real world based on visual input. Recent work [17] does sim2real transfer for audio-visual navigation by applying data augmentation empirically without measuring the acoustic gap. In this work, we systematically evaluate the sim2real gap by collecting data and identifying the spectral discrepancy.
@@ III-A SoundSpaces Platform and Audio-Visual Navigation
We first introduce the simulator that we are using. SoundSpaces 2.0 [19] is a state-of-the-art audio-visual simulation platform, which produces highly realistic audio and visual rendering for arbitrary camera and microphone locations in 3D scans of real-world environments [37, 43, 36]. It accounts for all major real-world acoustics phenomena: direct sounds, early specular/diffuse reflections, reverberation, spatialization, and materials and air absorption.
In audio-visual navigation, the goal is to navigate to a sounding object in an unknown environment by seeing and hearing. The location of the sound source is not known and needs to be inferred from the received audio. The agent needs to sample an action from {MOVE_FORWARD, TURN_LEFT, TURN_RIGHT, STOP} every step. If the agent issues the STOP action within a 1m radius of the goal location, the episode is considered successful.
There are different instantiations of the audio-visual navigation task, each with its goal specification. For example, AudioGoal [11] requires navigating to a static target in a discretized environment. Dynamic AudioGoal [44] requires navigating to a moving sound source. In this paper, we target the Continuous AudioGoal navigation benchmark introduced in SoundSpaces 2.0 [19] that relaxes the state space to be the continuous environment.
@@ III-B A Modular Design for Sim2real Transfer
The key challenge here is to formulate the proper waypoint prediction task that could survive the sim2real transfer. One existing approach [12] predicts the exact location of the audio goal directly, which is however an ill-posed problem since the environment geometry is unknown. For example, when the audio goal is in another room, the received audio reveals the direction to the door rather than the exact direction of the goal. We propose to predict the local acoustic field (sound pressure field) centered around the agent. Collecting acoustic field data for training is very expensive in the real world since it requires a complicated microphone array setup. However, rendering them in simulation is nearly free.
Figure 2: Acoustic field prediction model. The model first extracts audio and visual features, and then tiles, and concatenates both features to predict the acoustic field.
The hierarchical model alone however does not address the audio discrepancy directly. SoundSpaces 2.0 renders the audio propagation as a function of the geometry of the environment, the material properties, and source/receiver locations based on a bidirectional ray-tracing algorithm [45]. While it produces realistic audio renderings, there remains some difference between how sound propagates in the real world vs in simulation. It is known that ray-tracing-based algorithms yield worse performance with lower frequencies due to wave effects [20]. This implies the model needs to be aware of this spectral difference for sim2real. Thus we introduce a frequency-adaptive prediction strategy to help the model better transfer to the real world in Sec. III-E.
@@ III-C Acoustic Field Prediction
To tackle the acoustic field prediction problem, we first present a model that uses both audio and visual observations (see Fig. 2). The motivation behind using the visual sensor is that the visual observation of the surrounding environment can be useful in inferring the geometry of the environment, which affects the acoustic field. For example, walls often act as the boundary of the acoustic field.
More specifically, the model takes in a depth image of $128\times 128$ pixels. We use depth image instead of RGB image because it contains 3D information of the scene and tends to yield better performance [46, 47, 48]. The input audio to the model is a one-second segment of binaural audio, following prior work [11].
We use a pre-trained ResNet [49] to extract features from the input image and reshape it with a $1\times 1$ Conv layer into a 1d vector of size $512$. For the input binaural audio, we first process the waveforms with Short-Time Fourier Transform (STFT) to convert the time-domain signal into the frequency domain. We then use a 2D Conv net to encode the features and then tile and concatenate with the visual feature. Lastly, we feed the final output through one linear layer and reshape the prediction into the size of the target acoustic field $L\times L$.
@@ III-D2 Navigating to the Long-Term Goal
After sampling the long-term goal, for path planning, we use the Fast Marching Method (FMM) [50] to determine the best route to the goal in simulation. FMM takes the occupancy map, the agent’s current position, and the long-term goal as inputs. The occupancy map is computed by calculating the point cloud observed at each timestep using the depth camera. Next, FMM calculates the distance between each navigable point in the map to the long-term goal. The algorithm then selects the adjacent point on the map with the lowest value as its short-term goal and the agent then moves towards that point. When the long-term goal is sampled at a non-navigable location, we use breadth-first search (BFS) to find the closest available point to navigate to.
@@ III-E Frequency-adaptive Prediction
Existing audio-visual navigation models use all frequencies in the input audio. However, as discussed earlier, the acoustic gap with the real world is a function of frequencies. Thus models trained with all frequencies assuming them equally reliable would have lower performance when deployed on a real robot.
Given this observation, we first systematically examine how the gap changes as a function of the frequency. The idea is simple: with a given frequency band $[F_{1},F_{2}]$, we first train an acoustic field prediction (AFP) model in simulation using only that band, then test it on real-world data of the same sound and same band, and calculate the prediction error. We equally divide all frequencies into 5 subbands and we show the distribution of errors over the frequency bands in Fig. 4. As expected, the lower frequencies tend to yield larger prediction errors. However, the error does not monotonically decrease as frequency increases. We also trained a model that uses all frequencies, which has a distance error of $0.86m$, underperforming the best frequency band.
@@ IV-A SoundSpaces Acoustic Field Dataset
SoundSpaces 2.0 supports computing the impulse response $I(s,r)$ between the source location $s$ and the receiver location $r$ as a function of the 3D environment but does not have direct API support for rendering the acoustic field. To compute the field, given a $(s,r)$ pair, we first sample a grid centered at the receiver location of size $L\times L$, for each grid point $p$, we compute $I(s,p)$ which results in $L^{2}$ number of RIRs per receiver location. However, these RIRs are represented in the form of waveforms instead of single numbers. To best represent the sound pressure at each single point, we take the maximum amplitude of the waveform.
For sampling the S/R locations and environments, we utilize the existing dataset for audio-visual navigation, which provides configurations of the environment and source/receiver locations. This dataset uses scenes from the Matterport3D dataset [37], which contain scans of real-world environments such as apartments, offices, and even churches. We sample 500 episodes per environment for the 57 training environments in the navigation dataset. We also perform a similar operation to curate the validation and test set. In total, we collect 1.1M/52K/52K samples for train/val/test. Along with these acoustic fields, we also render the RGB-D images at the corresponding locations. See examples in Fig. 6.
@@ IV-B Real Measurements Collection
To measure the sim2real error, we collect real audio measurements to evaluate the trained model’s performance. For that, we use a 3Dio microphone to capture the binaural audio with a smartphone serving as the speaker output. We aligned the real-world parameters closely with those in our simulator, such as the height of the speaker and receiver. Since the simulator employs a mono receiver, the two-channel audio data we gathered is transformed into mono format by averaging the amplitude values across both channels. This process was repeated for ten distinct speaker positions (8 different directions w.r.t the agent and two data points for when the speaker is near the agent). We also downsample the acoustic field resolution from $9\times 9$ to $3\times 3$ so that we could collect more data in more environments.

### v2

@@ Abstract
Sim2real transfer has received increasing attention lately due to its success in transferring robotic policies learned in simulation to the real world. While significant progress has been made in transferring vision-based navigation policies, the current sim2real strategy for audio-visual navigation remains limited to basic data augmentation. Sound differs from light in that it spans across much wider frequencies and thus requires a different solution for sim2real. To understand how the acoustic sim2real gap varies with frequencies, we first define a novel acoustic field prediction (AFP) task that predicts the local sound pressure field. We then train frequency-specific AFP models in simulation and measure the prediction errors on collected real data. We propose a frequency-adaptive strategy that intelligently selects the best frequency band for prediction based on both the measured prior and the energy distribution of the received audio, which improves the generalization on real data. Coupled with waypoint navigation, we show the navigation policy not only improves navigation performance in simulation but also transfers successfully to real robots. This work demonstrates the potential of building autonomous agents that can see, hear, and act entirely from simulation, and transferring them to the real world.
@@ I Introduction
Navigation has been extensively studied in the robotics community and has been traditionally approached with Simultaneous Localization and Mapping (SLAM) [1] with Lidar sensors, which is limited in semantic reasoning. Recent research has increasingly focused on vision-centric navigation, where robots rely primarily on visual sensors to perceive their environment, showing significant success in photorealistic real-scanned settings [2]. Various tasks have been proposed, such as PointGoal navigation [3, 4, 5], ObjectGoal navigation  [6, 7], or visual exploration [8, 9, 10]. Other work further expands the sensory suite to include hearing. In particular, the AudioGoal task [11, 12] requires an agent to navigate to a sounding target (e.g., a ringing phone) using audio for directional and distance cues while using vision to avoid obstacles in the unmapped environment.
With the success of these learning-based navigation systems in simulation, efforts have been made to transfer learned policies to the real world by addressing the sim2real gap [13, 14, 15]. However, recent work [16] on audio-visual navigation with sim2real transfer through data augmentation, which fails to account for the influence the range of frequencies have on the sim2real gap. In this work, we systematically evaluate the acoustic gap and propose a solution to bridge it.
@@
State-of-the-art approaches in audio-visual navigation rely on reinforcement learning to train the navigation policy end-to-end [11, 17], which not only poses challenges for interpretation but also struggles to generalize to the real world due to various sim2real gaps. Recent advancements in visual navigation have shown successful sim2real transfer with hierarchical models [13, 8], which consist of a high-level path planner and a low-level motion planner. This hierarchical design helps mitigate some of the low-level physical discrepancies encountered during transfer. However, existing hierarchical models have not yet attempted to address sim2real for audio-visual navigation.
@@
Inspired by these methods, we design a modular approach to ease the transfer from simulation to the real world. To achieve this, we confront a key question: what is the proper high-level planning task that can survive sim2real transfer for audio-visual navigation? To this end, we propose a novel prediction task: acoustic field prediction—predicting the local sound pressure field around the agent. The gradient of this field reflects the direction of the sound. Measuring acoustic fields is expensive in the real world since it requires simultaneously capturing the sound pressure of all points in the field due to the dynamic nature of sound. However, computing acoustic fields is free in simulation. We first build an audio-visual model as the acoustic field predictor (AFP) and curate a large-scale acoustic field dataset on SoundSpaces 2.0 [18], the state-of-the-art audio-visual simulation platform. We show that this approach outperforms existing methods on the Continuous AudioGoal navigation benchmark.
@@
After validating the proposed approach in simulation, we then investigate where acoustic discrepancy arises. It is known that ray-tracing-based acoustic simulation algorithms introduce more errors with lower frequencies due to wave effects [19]. Given this observation, we focus on evaluating how sim2real errors vary with frequencies. We first collect real acoustic field data with the source sound being white noise, whose audio energy uniformly spans across all frequencies. We then train acoustic field prediction models that only take the sub-frequency band of the input audio and test it on the real white noise data. By computing the errors across multiple samples, we show that the errors do not strictly go down as the frequency goes higher, and using the best frequency band yields errors smaller than using all frequencies for the white noise sound.
@@
However, simply taking the best frequency band does not work for all sounds since different sounds have varying spectral distributions. To address this issue and make the model aware of the spectral difference, we propose a novel frequency-adaptive prediction strategy, which selects the optimal frequency sub-band based on measured errors and the spectral distribution of the received audio. To validate this approach, we collect more acoustic field data with various sounds and show that the frequency-adaptive model leads to the lowest error on the real data compared to other strategies.
@@
In summary, we propose a novel acoustic field prediction approach that learns to navigate without interaction with the environment. This approach improves the SOTA methods on the challenging Continuous AudioGoal navigation benchmark [18]. We perform a systematic evaluation of the sim2real challenge and propose a frequency-adaptive strategy as the treatment for sim2real. We show this strategy works on both collected real data as well as on our robot platform. To the best of our knowledge, this is the first work to investigate and propose a principled solution to the sim2real transfer problem for audio-visual navigation.
@@ II-A Embodied Navigation
To navigate autonomously, traditionally a robot builds a map via 3D reconstruction (i.e., SLAM) and then plans a path using the map [20]. Recent works have developed navigation policies that make navigation decisions in a previously unmapped environment from egocentric observations directly without relying on mapping [4, 7, 21]. Some recent efforts have developed audio-visual simulation platforms [11, 18, 22] that enable embodied agents to both see and hear. In the AudioGoal navigation task [11, 17], the agent must navigate to the source of sound in an unknown environment. The state-of-the-art audio-visual navigation models train policies with reinforcement learning (RL), requiring millions of samples. Inspired by recent work in learning potential functions for interaction-free navigation [8], we propose to predict acoustic fields for interaction-free audio-visual navigation.
@@ II-C Sim2real Transfer
Benefiting from recent large-scale datasets of real-world 3D scans [35, 36] and supporting simulators  [5, 37], recent works have shown success in enabling embodied agents in simulation. Transferring the model trained in simulation to the real world is thus of great interest. The mostly widely used approaches are domain randomization [38, 39], system identification [15, 40], and transfer learning and domain adaptation [41]. Most sim2real research studies transferring a policy from simulation to the real world based on visual input. Recent work [16] does sim2real transfer for audio-visual navigation by applying data augmentation empirically, which does not account for the effect of frequencies on sim2real.In this work, we systematically evaluate the sim2real gap by collecting data and identifying the spectral discrepancy.
@@ III-A SoundSpaces Platform and Audio-Visual Navigation
We first introduce the simulator that we are using. SoundSpaces 2.0 [18] is a state-of-the-art audio-visual simulation platform that produces highly realistic audio and visual rendering for arbitrary camera and microphone locations in 3D scans of real-world environments [36, 42, 35]. It accounts for all major real-world acoustics phenomena: direct sounds, early specular/diffuse reflections, reverberation, spatialization, and materials and air absorption.
In audio-visual navigation, the goal is to navigate to a sounding object in an unknown environment by seeing and hearing. The location of the sound source is not known and needs to be inferred from the received audio. At every step, the agent needs to sample an action from {MOVE_FORWARD, TURN_LEFT, TURN_RIGHT, STOP}. If the agent issues the STOP action within a 1m radius of the goal location, the episode is considered successful.
There are different instantiations of the audio-visual navigation task, each with its goal specification. For example, AudioGoal [11] requires navigating to a static target in a discretized environment, while Dynamic AudioGoal [43] requires navigating to a moving sound source. We target the Continuous AudioGoal navigation benchmark introduced in SoundSpaces 2.0 [18], which generalizes the state space to be the continuous environment.
@@ III-B A Modular Design for Sim2real Transfer
The key challenge here is to formulate the proper waypoint prediction task that could survive the sim2real transfer. One existing approach [12] predicts the exact location of the audio goal directly, which is however an ill-posed problem since the environment geometry is unknown. For example, when the audio goal is in another room, the received audio reveals the direction to the door rather than the exact direction of the goal. Instead of predicting the global audio goal location, we propose to predict the local acoustic field (sound pressure field) centered around the agent, which not only better captures the direction of the sound but also is more predictable from the visual observation of the environment.
Figure 2: Acoustic field prediction model. The model first extracts audio and visual features, and then tiles and concatenates features channel-wise to predict the acoustic field.
Defining the waypoint prediction task however does not address the audio discrepancy directly. SoundSpaces 2.0 renders the audio propagation as a function of the geometry of the environment, the material properties, and source/receiver locations based on a bidirectional ray-tracing algorithm [44]. While it produces realistic audio renderings, there remains some difference between how sound propagates in the real world vs in simulation. It is known that ray-tracing-based algorithms yield worse performance with lower frequencies due to wave effects [19]. This implies the model needs to be aware of this spectral difference for sim2real. Thus we introduce a frequency-adaptive prediction strategy to help the model better transfer to the real world in Sec. III-E.
@@ III-C Acoustic Field Prediction
We define acoustic field prediction as predicting the top-down sound pressure field of $L\times L$ centered at the agent, given the egocentric depth image of $128\times 128$ pixels and a one-second binaural audio.
To tackle the acoustic field prediction problem, we first present a model that uses both audio and visual observations (see Fig. 2). The motivation behind using the visual sensor is that the visual observation of the surrounding environment can be useful in inferring the geometry of the environment, which affects the acoustic field. For example, walls often act as the boundary of the acoustic field. More specifically, we use a pre-trained ResNet [45] to extract features from the input image and reshape it with a $1\times 1$ Conv layer into a 1d vector of size $512$. For the input binaural audio, we first process the waveforms with Short-Time Fourier Transform (STFT) to convert the time-domain signal into the frequency domain. We then use a 2D Conv net to encode the features and then tile and concatenate with the visual feature. Lastly, we feed the final output through one linear layer and reshape the prediction into the size of the target acoustic field $L\times L$.
@@ III-D2 Navigating to the Long-Term Goal
After sampling the long-term goal, for path planning, we use the Fast Marching Method (FMM) [46] to determine the best route to the goal in simulation. FMM takes the occupancy map that is built on the fly, the agent’s current position, and the long-term goal as inputs. The occupancy map is computed by calculating the point cloud observed at each timestep using the depth camera. Next, FMM calculates the distance between each navigable point in the map to the long-term goal. The algorithm then selects the adjacent point on the map with the lowest value as its short-term goal and the agent then moves towards that point. When the long-term goal is sampled at a non-navigable location, we use breadth-first search (BFS) to find the closest available point to navigate to.
@@ III-E Frequency-adaptive Prediction
Existing audio-visual navigation models use all frequencies in the input audio. However, sound spans across a wide range of frequencies, and due to imperfect geometric simulation techniques, the acoustic gap varies as a function of frequencies. Models trained with all frequencies assuming them equally reliable would have lower performance when deployed on a real robot.
Given this observation, we first systematically examine how the gap changes as a function of the frequency. The idea is simple: with a given frequency band $[F_{1},F_{2}]$, we first train an acoustic field prediction (AFP) model in simulation using only that band, then test it on real-world data of the same sound and same band, and calculate the prediction error. We equally divide all frequencies into $N$ subbands ($N=5$ based on hyperparameter tuning), and we show the distribution of errors over the frequency bands in Fig. 4, where error is defined as the distance between the predicted max and ground truth goal location. As expected, the lower frequencies tend to yield larger prediction errors. However, the error does not monotonically decrease as frequency increases likely due to simulation errors. We also trained a model that uses all frequencies, which has a distance error of $0.86m$, underperforming the best frequency band.
@@ IV-A SoundSpaces Acoustic Field Dataset
SoundSpaces 2.0 supports computing the impulse response $I(s,r)$ between the source location $s$ and the receiver location $r$ as a function of the 3D environment but does not have direct API support for rendering the acoustic field. To compute the field, given a $(s,r)$ pair, we first sample a grid centered at the receiver location of size $L\times L$, and for each grid point $p$, we compute $I(s,p)$ which results in $L^{2}$ number of RIRs per receiver location. However, these RIRs are represented in the form of waveforms instead of single numbers. To best represent the sound pressure at each single point, we take the maximum amplitude of the waveform.
For sampling the source/receiver locations and environments, we utilize the existing audio-visual navigation episodes [11], which provides configurations of the environment and source/receiver locations. This dataset uses scenes from the Matterport3D dataset [36], which contain scans of real-world environments such as apartments, offices, and even churches. We sample 500 episodes per environment for the 57 training environments in the navigation dataset. We also perform a similar operation to curate the validation and test set. In total, we collect 1.1M/52K/52K samples for train/val/test. Along with these acoustic fields, we also render the RGB-D images at the corresponding locations. See examples in Fig. 6.
@@ IV-B Real Measurements Collection
To measure the sim2real error, we collect real audio measurements to evaluate the trained model’s performance. For that, we use a 3Dio microphone to capture the binaural audio with a smartphone serving as the speaker output. We align the real-world parameters closely with those in our simulator, such as the height of the speaker and receiver. Since the simulator employs a mono receiver, the two-channel audio data we gather is transformed into mono format by averaging the amplitude values across both channels. This process is repeated for ten distinct speaker positions (8 different directions w.r.t the agent and two data points for when the speaker is near the agent). We also downsample the acoustic field resolution from $9\times 9$ to $3\times 3$ so that we can collect more data in more environments.